element X reacts with oxygen to produce a pure sample of X2O3 . in an experiment it is found that 1.000g of X produce 1.1596g of X2O3 .using the known atomic wieght of oxygen 15.9994g/mol calculate atomic wieght of X

Grade:12

3 Answers

Arun
6 years ago
Dear student
 
the reaction equation is 2X + 3O ----> X2O3 2 moles X produce 1 mole X2O3 let the molecular mass of X = M, so 1 g X = 1/M moles X, which produce 1/(2M) moles X2O3 ---> 1.1596 g X2O3 = 1/(2M) moles the molecular mass of X2O3 is (2M+ 3*16) = (2M+48) so 1.1596 g are 1.1596/(2M+48) moles together it follows, that 1/(2M) moles = 1.1596/(2M + 48) moles 2M+48 = 1.1596*2M 48 = 0.3192M M = 150.376 g
Regards
Arun (askIITians forum ezpere)
Srishti
5 years ago
X + O2 → X2O3 ..
X is 1 gm X2O3 is 1.1596 gm so O2 is 0.1596 gm 
0.1596 gm react with 1 gm X2 and 8 gm O2 react with 8/0.1596 gm X2 which is the equivalent weight of X2 ...so according to formula E =M/n where n is the n factor of X which is 3 in this rxn.. We get 8/0.1596 = M/3 ..and We get M =150.4
Abd monther hamdan
2 years ago
Element X react with oxygen with a formula X₂O3 what is the correct formula when the same element react with floure (F)